# A Closer Look at Time Steps is Worthy of Triple Speed-Up for Diffusion Model Training

### v2

@@ Abstract
Training diffusion models is always a computation-intensive task. In this paper, we introduce a novel speed-up method for diffusion model training, called SpeeD, which is based on a closer look at time steps. Our key findings are: i) Time steps can be empirically divided into acceleration, deceleration, and convergence areas based on the process increment. ii) These time steps are imbalanced, with many concentrated in the convergence area. iii) The concentrated steps provide limited benefits for diffusion training. To address this, we design an asymmetric sampling strategy that reduces the frequency of steps from the convergence area while increasing the sampling probability for steps from other areas. Additionally, we propose a weighting strategy to emphasize the importance of time steps with rapid-change process increments. As a plug-and-play and architecture-agnostic approach, SpeeD consistently achieves 3$\times$ acceleration across various diffusion architectures, datasets, and tasks. Notably, due to its simple design, our approach significantly reduces the cost of diffusion model training with minimal overhead. Our research enables more researchers to train diffusion models at a lower cost.
@@ 1 Introduction
Recently, some acceleration methods for diffusion training focus on time steps, primarily using re-weighting and re-sampling 1) Re-weighting on the time steps based on heuristic rules. P2 (Choi et al., 2022) and Min-SNR (Hang et al., 2023) use monotonous and single-peak weighting strategies according to sign-to-noise ratios (SNR) in different time steps. 2) Re-sampling the time steps. Log-Normal (Karras et al., 2022) assigns high sampling probabilities for the middle time steps of the diffusion process. CLTS (Xu et al., 2024) proposes a curriculum learning based time step schedule, gradually tuning the sampling probability from uniform to Gaussian by interpolation for acceleration as shown in Fig. 1(b).
To investigate the essence of the above accelerations, we take a closer look at the time steps. The diffusion models essentially learn and estimate the process increment noted as $\delta_{t}:=x_{t+1}-x_{t}$ at time step $t$ at training and inference phase. As illustrated in the left of Fig. 1(a), we visualize the changes of mean and variance of process increment through the time steps. The time steps are categorized into three areas: acceleration, deceleration, and convergence, based on observations and theoretical analyses (Sec. 2.2) that both the changes of mean and variance initially accelerate, subsequently decelerate, and ultimately converge in a narrow interval.
One can easily find that the proportions of the three areas are imbalanced: there are a large number of time steps at convergence area, and the others are small. Another finding is that process increments at the convergence area are almost identical noise, e.g., in DDPM, the distribution are nearly $\mathcal{N}(0,2\mathbf{I})$, where $\mathbf{I}$ is the unit matrix. To further explore the characteristics of these three areas, we visualize the training loss curve in the right of Fig. 1(a). The loss values from the convergence area are much lower than those from the other areas, which indicates estimating the identical noise is easy.
Previous acceleration works have achieved promising results, but the analysis of time steps remains relatively under-explored. P2 (Choi et al., 2022) and Min-SNR (Hang et al., 2023) are two re-weighting methods, with their weighting curves across time steps as shown in Fig. 1(b). They employ uniform sampling of time steps, which include too many easy samples from the convergence area during diffusion model training.
On the other hand, most re-sampling methods heuristically emphasize sampling the middle-time steps, but they do not dive into the difference between the acceleration and convergence areas. For example, CLTS (Xu et al., 2024) gradually changes the sampling distribution from uniform to Gaussian by interpolation as shown in Fig. 1(b). The sampling probability of the acceleration area drops faster than the one of the convergence area. The acceleration area is still under-sampled and therefore not well-learned.
@@
Figure 1: Motivation: designing an efficient training via analyzing process increment $\delta_{t}$ at different time steps. (a) $\mathbf{E}(\delta_{t})$ and $\text{Var}(\delta_{t})$ are the mean and variance of process increments $\delta_{t}$. Two histograms represent the proportions of the process increments at different noise levels (left) and the proportions of the time steps (right) in the three areas. The loss curve is obtained from DDPM (Ho et al., 2020) on CIFAR-10 (Krizhevsky et al., 2009). (b) $w(t)$ and $\mathbf{P}(t)$ are respectively the weighting and sampling curve. The probability of convergence area being sampled remains, while the one of acceleration is reduced faster.
@@
Our approach has the following characteristics: SpeeD is compatible with various diffusion model training methods, i.e., U-Net (Ronneberger et al., 2015) and DiT (Peebles & Xie, 2023) with minimal modifications. For performance, SpeeD achieves non-trivial improvements than baseline and other methods at the same training iterations. For efficiency, SpeeD consistently accelerates the diffusion training by 3$\times$ across various tasks and datasets. It helps mitigate the heavy computational cost for diffusion model training, enabling more researchers to train a model at an acceptable expense. The extra time complexity of SpeeD is $\mathcal{O}(1)$, costing only seconds to reduce days of diffusion models training on datasets like FFHQ, MetFaces and ImageNet-1K. We hope this work can bring novel insights for efficient diffusion model training.
2 Speeding Up Training from Process Increments at Time Steps
@@ 2.1 Preliminaries of Diffusion Models
In conventional DDPM (Ho et al., 2020; Sohl-Dickstein et al., 2015), given data $x_{0}\sim p(x_{0})$, the forward process is a Markov-Gaussian process that gradually adds noise to obtain a perturbed sequence {$x_{1},x_{2},...,x_{T}$},
$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathrm{\textbf{I}}),q(x_{1:T}|x_{0})=\prod_{t=1}^{T}q(x_{t}|x_{t-1}),$
where $\mathbf{I}$ is the unit matrix, $T$ is the total number of time steps, $q$ and $\mathcal{N}$ represent forward process and Gaussian distribution parameterized by scheduled hyper-parameters $\{\beta_{t}\}_{t\in[T]}$. Perturbed samples are sampled by $x_{t}=\sqrt{\bar{\alpha}_{t}}\cdot x_{0}+\sqrt{1-\bar{\alpha}_{t}}\cdot\epsilon,\epsilon\sim\mathcal{N}(0,\mathbf{I})$, where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$.
For diffusion model training, the forward process is divided into pairs of samples and targeted process increments by time steps $t$, defined as $\delta_{t}:=x_{t+1}-x_{t}$. The diffusion model is expected to predict the next step from the given time step. The training loss (Ho et al., 2020) for diffusion models is to predict the normalized noise. The loss highlighted with weighting and sampling modules:
$L=\mathbb{E}_{\mu_{t}}[w_{t}||\epsilon-\epsilon_{\theta}(x_{t},t)||^{2}]:=\int_{t}w_{t}||\epsilon-\epsilon_{\theta}(x_{t},t)||^{2}\mathbf{d}\mu_{t},\epsilon\sim\mathcal{N}(0,\mathbf{I}).$ (1)
@@
2.2 Case Study on DDPM: A Closer Look at Time Steps
In DDPM, the diffusion model learns the noise added in the forward process at given $t^{\text{th}}$ time step. The noise is presented as $\epsilon$, the label in Eqn. 1, which is the normalized process increment at given time step. This label tells what the output of the diffusion model is aligning to. To take a closer look, we focus on the nature of the process increment $\delta_{t}$ itself to study the diffusion process $x_{t}\rightarrow x_{t+1}$, instead of $\epsilon$ the normalized one. According to Theorem 1 and Remark 1, based on the variation trends of process increments $\delta_{t}$, we can distinguish three distinct areas: acceleration, deceleration, and convergence. The detailed discussion is shown as follows:
@@ 2.2 Case Study on DDPM: A Closer Look at Time Steps
In DDPM’s setting (Ho et al., 2020), the linear schedule hyper-parameters $\{\beta_{t}\}_{t\in[T]}$ is an equivariant series, the extreme deviation $\Delta_{\beta}:=\max_{t}\beta_{t}-\min_{t}\beta_{t}$, $T$ is the total number of time steps, and we have the bounds about the process increment $\delta_{t}\sim\mathcal{N}(\phi_{t},\Psi_{t})$, where $\phi_{t}:=(\sqrt{\alpha_{t+1}}-1)\sqrt{\bar{\alpha}_{t}}x_{0},\Psi_{t}:=[2-\bar{\alpha}_{t}(1+\alpha_{t+1})]\mathbf{I}$, $\mathbf{I}$ is the unit matrix:
Upper-bound: $||\phi_{t}||^{2}$ $\leq\hat{\phi}_{t}||\mathbb{E}x_{0}||^{2},$ $\hat{\phi}_{t}:=\beta_{\max}\exp\{-(\beta_{0}+{\Delta_{\beta}t}/{2T})t\}$ (2)
Lower-bound: $\Psi_{t}$ $\succeq\hat{\Psi}_{t}\mathbf{I},$ $\hat{\Psi}_{t}:=2-2\exp\{-(\beta_{0}+{\Delta_{\beta}t}/{2T})t\}$
@@
In the convergence area, the variations of $\delta_{t}$ stabilize, signifying that the process is approaching a steady state.
@@
2.3 General Cases Beyond DDPM: Dive into The Commonality.
This section generalize the above findings on DDPM to broader settings. The findings are about the process increments $\delta_{t}:=x_{t+1}-x_{t}$, and the related differentiation of right limit $\mathbf{d}x$ in forward process. The corresponding SDE (Karras et al., 2022) and the discretization are:
$\mathbf{d}x=x\dot{s}/s\mathbf{d}t+s\sqrt{\dot{\sigma}\sigma}\mathbf{d}w,\qquad x_{t}=s_{t}x_{0}+s_{t}\sigma_{t}\epsilon,\epsilon\sim\mathcal{N}(0,\mathbf{I}),$
where scale factor $s=s_{t}$ and noise standard deviation (std.) $\sigma=\sigma_{t}$ are the main designs related to the main factors $\hat{\phi}_{t}$ and $\hat{\Psi}_{t}$ about process increment $\delta_{t}$ in Theorem 1 across time steps $t$.
Generalize Theorem 1: $s$-$\sigma$ Scheduled Process Increments.
The generalized process increment is $\delta\sim\mathcal{N}(\Delta x_{0},\Sigma\mathbf{I})$, where $\Delta:=s_{+}-s$ and $\Sigma:=s_{+}^{2}\sigma_{+}^{2}+s^{2}\sigma^{2}$ across $t$. $\Delta$, $\Sigma$ are continuous on $t$ without discretization, where $s_{+}$ and $\sigma_{+}$ are the right outer limits, i.e., $s(t+\mathbf{d}t)$. In discretization, $\Delta_{t}$ and $\Sigma_{t}$, marked by $t$, are related to sample granularity of time step $t$. Like Theorem 1, we study the variation of process increments by $\dot{\Delta}=\dot{s}_{+}-\dot{s}$ and $\dot{\Sigma}=\mathbf{m}^{\top}\dot{\mathbf{n}}$ where $\mathbf{m}=[s_{+}^{2},\sigma_{+}^{2},s^{2},\sigma^{2}]^{\top}$, $\mathbf{n}=[\sigma_{+}^{2},s_{+}^{2},\sigma^{2},s^{2}]^{\top}$. This formulation involves only terms about derivatives of given schedule functions, which brings computational convenience. Tab. 1 provides all ingredients needed to calculate curves of $\dot{\Delta}$ and $\dot{\Sigma}$ in schedules of VP, VE and EDM.
We also generalize the previous takeaways from the DDPM case study to $s$-$\sigma$ scheduled setting with the following analyses and discussion.
@@
Sampling deserves more attention. The SDE design goal of most diffusion models nowadays is to add much larger noise to the data so that the samples can cover larger space. However, in terms of the process increment, it always results in a low signal-to-noise ratio of the late data when $t$ is large. Either the standard deviation is too large or the $s$ used to suppress it is too small. For instance, EDM does not bias the data distribution from expectations due to the scale $\dot{\Delta}=0$, but $\dot{\Sigma}=2[(t+\mathbf{d}t)^{2}+t^{2}]$ is a quadratic increase as $t$ grows. In VP, $s\rightarrow 0$, as $t$ grows, leads that the model needs to recover expectations in approximation. For these samples, which are not very informative, a single weight adjustment is not as efficient as directly reducing the sampling rate.
Table 1: The ingredients of generalized curves $\dot{\Delta}$ and $\dot{\Sigma}$ schedules about mainstream SDE designs, including VP, VE (Song et al., 2021), EDM (Karras et al., 2022).
Schedules $s$ $\sigma^{2}$ $\dot{s}$ $\dot{\sigma}$
VE $1$ $t$ 00 $1$
EDM $1$ $t^{2}$ 00 $2t$
2.4 Overview of SpeeD
Based on the above observations and analyses, we propose SpeeD , a novel approach for achieving lossless training acceleration tailored for diffusion models. As illustrated in Fig. 2, SpeeD suppresses the trivial time steps from convergence area, and weight the rapid-change intervals between acceleration and deceleration areas. Correspondingly, two main modules, asymmetric sampling and change-aware weighting, are proposed. Asymmetric sampling uses a two-step step function to respectively suppress and increase the sampling probability corresponding to trivial and beneficial time steps. Change-aware weighting is based on the change rate of process increment $\partial_{t}\Psi(t)$.
2.5 Asymmetric Sampling
SpeeD adopts the time steps sampling probability $\mathbf{P}(t)$ as the step function in Eqn. 3 to construct the loss in Eqn. 1. We first define $\tau$ as the step threshold in $\mathbf{P}(t)$. The pre-defined boundary $\tau$ means the area where the time step are suppressed. The sampling probability is $k$ times from time-steps where $t<\tau$ than that from $t>\tau$ instead of the uniform sampling $\mathbf{U}(t)=1/T$.
$\mathbf{P}(t)=\begin{cases}\frac{k}{T+\tau(k-1)},\quad&0<t\leq\tau\\ \frac{1}{T+\tau(k-1)},\quad&\tau<t\leq T\end{cases},\quad\text{where suppression intensity }k\geq 1\text{ and }\tau\in(0,T].$ (3)
Threshold Selection $\tau$.
2.6 Change-Aware Weighting

### v3

@@ Abstract
Training diffusion models is always a computation-intensive task. In this paper, we introduce a novel speed-up method for diffusion model training, called SpeeD, which is based on a closer look at time steps. Our key findings are: i) Time steps can be empirically divided into acceleration, deceleration, and convergence areas based on the process increment. ii) These time steps are imbalanced, with many concentrated in the convergence area. iii) The concentrated steps provide limited benefits for diffusion training. To address this, we design an asymmetric sampling strategy that reduces the frequency of steps from the convergence area while increasing the sampling probability in other areas. Additionally, we propose a weighting strategy to emphasize the importance of time steps with rapid-change process increments. As a plug-and-play and architecture-agnostic approach, SpeeD consistently achieves 3$\times$ acceleration across various diffusion architectures, datasets, and tasks. Notably, due to its simple design, our approach significantly reduces the cost of diffusion model training with minimal overhead. Our research enables more researchers to train diffusion models at a lower cost.
00footnotetext: equal contribution, corresponding author.
@@ 1 Introduction
Recently, some acceleration methods for diffusion training focus on time steps, primarily using re-weighting and re-sampling 1) Re-weighting on the time steps based on heuristic rules. P2 [8] and Min-SNR [17] use monotonous and single-peak weighting strategies according to sign-to-noise ratios (SNR) in different time steps. 2) Re-sampling the time steps. Log-Normal [28] assigns high sampling probabilities for the middle time steps of the diffusion process. CLTS [67] proposes a curriculum learning based time step schedule, gradually tuning the sampling probability from uniform to Gaussian by interpolation for acceleration as shown in Fig. 2.
To investigate the essence of the above accelerations, we take a closer look at the time steps.
@@
Figure 1: Closer look at time steps: More than half of the time steps are almost pure noise and easy-to-learn. Motivation: designing an efficient training via analyzing process increment $\delta_{t}$ at different time steps. $\mathbf{E}(\delta_{t})$ and $\text{Var}(\delta_{t})$ are the mean and variance of process increments $\delta_{t}$. Two histograms represent the proportions of the process increments at different noise levels (left) and the proportions of the time steps (right) in the three areas. The loss curve is obtained from DDPM [20] on CIFAR-10 [33].
As shown in the left of Fig. 1, we visualize the changes of mean and variance of the process increment $\delta_{t}:=x_{t+1}-x_{t}$ at time step $t$. The time steps are divided into three areas: acceleration, deceleration, and convergence. We identify the characteristics of three areas of the time steps. One can easily find that the proportions of the three areas are imbalanced: a much larger number of time steps at the narrow convergence area. Besides, process increments at the convergence area are almost identical noise, e.g., in DDPM, the distribution are almost pure white noise. To further explore the characteristics of these three areas, we visualize the training loss curve in the right of Fig. 1. The loss values from the convergence area are much lower than the others, which indicates estimating the identical noise is easy.
Figure 2: Re-weighting and re-sampling methods can’t eliminate the redundancy and under-sample issues. $w(t)$ and $\mathbf{P}(t)$ are respectively the weighting and sampling curve. The probability of convergence area being sampled remains, while the one of acceleration is reduced faster.
Previous acceleration works have achieved promising results, but the analysis of time steps remains relatively under-explored. P2 [8] and Min-SNR [17] are two re-weighting methods, with their weighting curves across time steps as shown in Fig. 2. They employ uniform sampling of time steps, which include too many easy samples from the convergence area during diffusion model training. Most re-sampling methods heuristically emphasize sampling the middle-time steps, but they do not dive into the difference between the acceleration and convergence areas. CLTS [67] gradually changes the sampling distribution from uniform to Gaussian by interpolation as shown in Fig. 2. The sampling probability of the acceleration area drops faster than the one of the convergence area. The acceleration area is still under-sampled and therefore not well-learned.
@@
Our approach has the following characteristics: SpeeD is compatible with various diffusion model training methods, i.e., U-Net [52] and DiT [47] with minimal modifications. For performance, SpeeD achieves non-trivial improvements than baseline and other methods at the same training iterations. For efficiency, SpeeD consistently accelerates the diffusion training by 3$\times$ across various tasks and datasets. It helps mitigate the heavy computational cost for diffusion model training, enabling more researchers to train a model at an acceptable expense. The extra time complexity of SpeeD is $\mathcal{O}(1)$, costing only seconds to reduce days of diffusion models training on datasets like FFHQ, MetFaces and ImageNet-1K. We hope this work can bring novel insights for efficient diffusion model training.
2 Speeding Up Training: Time Steps
@@ 2.1 Preliminaries of Diffusion Models
In conventional DDPM [20, 57], given data $x_{0}\sim p(x_{0})$, the forward process is a Markov-Gaussian process that gradually adds noise to obtain a sequence {$x_{1},x_{2},...,x_{T}$},
$q(x_{t}|x_{t-1})$ $=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathrm{\textbf{I}}),$
$q(x_{1:T}|x_{0})$ $=\prod_{t=1}^{T}q(x_{t}|x_{t-1}),$
where $\mathbf{I}$ is the unit matrix, $T$ is the total number of time steps, $q$ and $\mathcal{N}$ represent forward process and Gaussian dist. parameterized by scheduled hyper-parameters $\{\beta_{t}\}_{t\in[T]}$. Perturbed samples are sampled by $x_{t}=\sqrt{\bar{\alpha}_{t}}\cdot x_{0}+\sqrt{1-\bar{\alpha}_{t}}\cdot\epsilon,\epsilon\sim\mathcal{N}(0,\mathbf{I})$, where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$.
For diffusion model training, the forward process is divided into pairs of samples and targeted process increments by time steps $t$, defined as $\delta_{t}:=x_{t+1}-x_{t}$. The diffusion model is expected to predict the next step from the given time step. The training loss [20] for diffusion models is to predict the normalized noise. The loss highlighted with weighting and sampling modules, where $\epsilon\sim\mathcal{N}(0,\mathbf{I})$:
$L=\mathbb{E}_{\mu_{t}}[w_{t}||\epsilon-\epsilon_{\theta}(x_{t},t)||^{2}]:=\int_{t}w_{t}||\epsilon-\epsilon_{\theta}(x_{t},t)||^{2}\mathbf{d}\mu_{t},$ (1)
@@
2.2 Overview of SpeeD
Based on the above observations and analyses, we propose SpeeD , a novel approach for achieving lossless training acceleration tailored for diffusion models. As illustrated in Fig. 3, SpeeD suppresses the trivial time steps from convergence area, and weight the rapid-change intervals between acceleration and deceleration areas. Correspondingly, two main modules, asymmetric sampling and change-aware weighting, are proposed. Asymmetric sampling uses a two-step step function to respectively suppress and increase the sampling probability corresponding to trivial and beneficial time steps. Change-aware weighting is based on the change rate of process increment $\partial_{t}\Psi(t)$ in Theorem 1.
2.3 Asymmetric Sampling
SpeeD adopts the time steps sampling probability $\mathbf{P}(t)$ as the step function in Eqn. 2 to construct the loss in Eqn. 1. We first define $\tau$ as the step threshold in $\mathbf{P}(t)$. The pre-defined boundary $\tau$ means the area where the time step are suppressed. The sampling probability is $k$ times from time-steps where $t<\tau$ than that from $t>\tau$ instead of the uniform sampling $\mathbf{U}(t)=1/T$.
$\mathbf{P}(t)=\begin{cases}\frac{k}{T+\tau(k-1)},\quad&0<t\leq\tau;\\ \frac{1}{T+\tau(k-1)},\quad&\tau<t\leq T,\end{cases}$ (2)
where suppression intensity $k\geq 1\text{ and }\tau\in(0,T]$.
Threshold Selection $\tau$.
2.4 Change-Aware Weighting
2.5 Case Study: DDPM
In DDPM, the diffusion model learns the noise added in the forward process at given $t^{\text{th}}$ time step. The noise is presented as $\epsilon$, the label in Eqn. 1, which is the normalized process increment at given time step. This label tells what the output of the diffusion model is aligning to. To take a closer look, we focus on the nature of the process increment $\delta_{t}$ itself to study the diffusion process $x_{t}\rightarrow x_{t+1}$, instead of $\epsilon$ the normalized one. According to Theorem 1 and Remark 1, based on the variation trends of process increments $\delta_{t}$, we can distinguish three distinct areas: acceleration, deceleration, and convergence.
@@ 2.5 Case Study: DDPM
In DDPM’s setting [20], the linear schedule hyper-parameters $\{\beta_{t}\}_{t\in[T]}$ is an equivariant series, the extreme deviation $\Delta_{\beta}:=\max_{t}\beta_{t}-\min_{t}\beta_{t}$, $T$ is the total number of time steps, and we have the bounds about the process increment $\delta_{t}\sim\mathcal{N}(\phi_{t},\Psi_{t})$, where $\phi_{t}:=(\sqrt{\alpha_{t+1}}-1)\sqrt{\bar{\alpha}_{t}}x_{0},\Psi_{t}:=[2-\bar{\alpha}_{t}(1+\alpha_{t+1})]\mathbf{I}$, $\mathbf{I}$ is the unit matrix:
[…]
where ${}\hat{\phi}_{t}:=\beta_{\max}\exp\{-(\beta_{0}+{\Delta_{\beta}t}/{2T})t\}$ and ${}\hat{\Psi}_{t}:=2-2\exp\{-(\beta_{0}+{\Delta_{\beta}t}/{2T})t\}$.
@@
In the convergence area, the variations of $\delta_{t}$ stabilize, signifying that the process is getting steady.
@@
2.6 General Cases: Beyond DDPM
This section generalize the above findings on DDPM to broader settings. The findings are about the process increments $\delta_{t}:=x_{t+1}-x_{t}$, and the related differentiation of right limit $\mathbf{d}x$ in forward process. The corresponding SDE [28] and the discretization are:
$\mathbf{d}x=x\dot{s}/s\mathbf{d}t+s\sqrt{\dot{\sigma}\sigma}\mathbf{d}w,~{}x_{t}=s_{t}x_{0}+s_{t}\sigma_{t}\epsilon,\epsilon\sim\mathcal{N}(0,\mathbf{I}),$
where scale factor $s=s_{t}$ and noise standard deviation (std.) $\sigma=\sigma_{t}$ are the main designs related to the factors $\hat{\phi}_{t}$ and $\hat{\Psi}_{t}$ about process increment $\delta_{t}$ in Theorem 1 at time steps $t$.
Generalize Theorem 1
: $s$-$\sigma$ Scheduled Process Increments. The generalized process increment is $\delta\sim\mathcal{N}(\Delta x_{0},\Sigma\mathbf{I})$, where $\Delta:=s_{+}-s$ and $\Sigma:=s_{+}^{2}\sigma_{+}^{2}+s^{2}\sigma^{2}$ across $t$. $\Delta$, $\Sigma$ are continuous on $t$ without discretization, where $s_{+}$ and $\sigma_{+}$ are the right outer limits, i.e., $s(t+\mathbf{d}t)$. In discretization, $\Delta_{t}$ and $\Sigma_{t}$, marked by $t$, are related to sample granularity of time step $t$. Like Theorem 1, we study the variation of process increments by $\dot{\Delta}=\dot{s}_{+}-\dot{s}$ and $\dot{\Sigma}=\mathbf{m}^{\top}\dot{\mathbf{n}}$ where $\mathbf{m}=[s_{+}^{2},\sigma_{+}^{2},s^{2},\sigma^{2}]^{\top}$, $\mathbf{n}=[\sigma_{+}^{2},s_{+}^{2},\sigma^{2},s^{2}]^{\top}$. This formulation involves only terms about derivatives of given schedule functions, which brings computational convenience. Tab. 8 (in Appendix) provides all ingredients needed to calculate curves of $\dot{\Delta}$ and $\dot{\Sigma}$ in schedules of VP, VE and EDM. We also generalize the previous takeaways from the DDPM case study to $s$-$\sigma$ scheduled setting with the following analyses.
@@ 2.6 General Cases: Beyond DDPM
Sampling deserves more attention. The SDE design goal of most diffusion models nowadays is to add much larger noise to the data so that the samples can cover larger space. However, in terms of the process increment, it always results in a low signal-to-noise ratio of the late data when $t$ is large. Either the standard deviation is too large or the $s$ used to suppress it is too small. For instance, EDM does not bias the data distribution from expectations due to the scale $\dot{\Delta}=0$, but $\dot{\Sigma}=2[(t+\mathbf{d}t)^{2}+t^{2}]$ is a quadratic increase as $t$ grows. In VP, $s\rightarrow 0$, as $t$ grows, leads that the model needs to recover expectations in approximation. For these samples, which are not very informative, a single weight adjustment is not as efficient as reducing the sampling rate.